OpenStudy (zenmo):

Help with factoring (2+h)^3

OpenStudy (zenmo):

\[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (zenmo):

\[(2+h)(4-2h^2+h^2)=8-4h+2h^2+4h-2h^2+h^3\] is what I get.

OpenStudy (zenmo):

\[8+h^3\]

OpenStudy (lauren32583):

Make sure that you put your answer in decending order like so h^3+6h^2+12h+8

OpenStudy (zenmo):

How to get the answer?

OpenStudy (zenmo):

I only get \[8+h^3\] when i factor out \[(2+h)(4-2h+h^2)\] from \[(2+h)\]

OpenStudy (zenmo):

\[(2+h)^3\]

OpenStudy (firekat97):

You shouldn't be using that rule, that rule is for \[a^3 + b^3 = (a + b) ( a^2 - ab + b^2)\]

OpenStudy (lauren32583):

you multiply (2+h) times itself 3 times bc it is cubed (2+h)(2+h)(2+h) *I distribute the first 2 then I distribute to the 3rd one. 4+ 2h+2h+h^2(2+h)

OpenStudy (firekat97):

what you want to do is simplify \[(2 + h) (2 + h)^2\]

OpenStudy (lauren32583):

then all you do is distribute and combine like terms.

OpenStudy (zenmo):

Oh, I got it now. Thanks all.